OpenStudy (anonymous):

What is the value of the expression 1 ----- -3 4

OpenStudy (anonymous):

1/(4)^-3 ???

OpenStudy (anonymous):



OpenStudy (anonymous):

4^-3 = 1/64 = 1/(1/64) Are you with me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you have a fraction over a fraction, you can reverse it through multiplication 2/(3/4) = 4(2)/3 Now to your problem: 1/(1/64) = 64(1)/1

OpenStudy (anonymous):